OpenStudy (anonymous):

Decompose (x - 1)/(3x^2 - 14 x + 15) into partial fractions

OpenStudy (anonymous):

2. (x - 1)/((3x - 5)(x - 3)) x - 1 = A/(3x - 5) + B/(x - 3) x - 1 = A(x - 3) + B(3x - 5) x - 1 = Ax - 3A + 3Bx - 5B x - 1 = x(A + 3B) - 3A - 5B 3(A + 3B = 1) -3A - 5B = -1 3A + 9B = 3 -3A - 5B = -1 4B = 2 B = ½ A + 3(½) = 1 A = -½

OpenStudy (anonymous):

can u help me too

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

your wellcome hold on let me copy it